Problem 6-05A a1-a3, b (Part Level Submission) (Video)

You are provided with the following information for Larkspur Inc. for the month ended June 30, 2020. Larkspur uses the periodic method for inventory.

Date

Description

Quantity

Unit Cost or
Selling Price

June 1 Beginning inventory 45 $42
June 4 Purchase 136 46
June 10 Sale 113 69
June 11 Sale return 15 69
June 18 Purchase 54 48
June 18 Purchase return 12 48
June 25 Sale 68 74
June 28 Purchase 35 52
Weighted-average cost per unit $

46.44

Calculate ending inventory, cost of goods sold, gross profit under each of the following methods. (1) LIFO. (2) FIFO. (3) Average-cost. (Round average-cost method answers to 2 decimal places, e.g. 1,250.25 and other answers to 0 decimal places, e.g. 1,250. Use weighted-average unit cost rounded to 2 decimal places for computations.)

LIFO

FIFO

AVERAGE-COST

The ending inventory $ $ $
The cost of goods sold $ $ $
Gross profit $ $ $
Answer #1

The answer has been presented in the supporting sheet. For detailed answer refer to the supporting sheet.

Answer Firstly we have to calculate the cost of goods available for sale = No. of units * purchase price) = (45*42) + (136*46

A B D E F 7 3 Ending Inventory LIFO 9 = no. of units in ending inventory * Cost as per FIFO 0 = (45*42) + ((92-45) *46) 1 = $